OpenStudy (anonymous):

Show all steps and give your answer correct to 3 decimal places. Solve for x in the equation 3(x-2) = 4

OpenStudy (anonymous):

3(x-2) = 4 3x - 6 = 4 3x = 4 + 6 3x = 10 x = (10/3) = 3.3333333333333333333333333333333333